Ella’s grandmother sewed her change purse. When she gifted it to her, Ella counted 23 coins already in it Ella noticed there were only dimes and quarters worth $4.70 how many dimes for in the change purse
Answer by ikleyn(52797)

.

x quarters and (23-x) dimes.


Total money equation


    10*(23-x) + 25x = 470  cents

    10*23 - 10x + 25x = 470

    25x - 10x = 470 - 10*23

    15x       = 240

      x       = 240/15 = 16.


ANSWER.  16 quarters and the rest, 23-16 = 7 dimes.

Solved.